Which one of the following will increase the current ratio but not the quick ratio? O increase in inventory O decrease in cash O increase in accounts payable decrease in accounts receivable
Welcome Inn has total equity of $471.000 and a debt-equity ratio of .54. What is the firms equity multiplier? O 1.54 O 1.40 ○ .46 O 185
The debt-equity ratio is equal to which one of the following? O Equity multiplier + 1 O Long-term debt / Total equity O Equity multiplier- 1 O Long-term debt + Total equity) /Total equity
Arlenes Outlet has sales of $384,000. Costs of goods sold equal 64 percent of sales. The store has $39,250 in inventory. On average, how long does inventory set on the shelf before it is sold? O 37.31 days O 58.29 days 65.37 days O3871 days
Answer #1

Part 1

Increase in Inventory

Increase in Stock will result in increase in current asset hence current asset ratio will increase

Part 2

Equity multiplier = Total assets / Total equity

Total asset =Debt + Equity

Debt/Equity =471000*.54

Equity multiplier =(471000*.54+471000)/471000=1.54

Part 3

Debt equity ratio =Debt/Equity

Part 4

Inventory turnover ratio =Cost of goods sold/Average inventory

=($384000*.64)/$39250=6.27 times

How long inventory set on the shelf=365/Inventory turnover ratio =365/6.27 =58.29 days